# Explicit Second-Order Min-Max Optimization Methods with Optimal Convergence Guarantee

### v4

@@ 4.3 Inexact subproblem solving and complexity analysis
We clarify how to obtain an inexact solution of the nonlinear equation problem in Eq. (4.1) such that Condition 4.2 holds. To that end, we present a new subroutine that solves each subproblem using a single Schur decomposition and $O(\log\log(1/\epsilon))$ calls to a linear system solver in a quasi-upper-triangular system. This gives a total complexity of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ which outperforms that of $O((m+n)^{\omega}\epsilon^{-2/3}\log\log(1/\epsilon))$ achieved by the best known line-search-based method.
For ease of presentation, we omit the subscript and rewrite the nonlinear equation problem in Eq. (4.1) as
@@
Although $J(\hat{\mathbf{z}})$ is not symmetric, it has a Schur decomposition $J(\hat{\mathbf{z}})=QUQ^{-1}$ where $U$ is quasi-upper-triangular (since all entries of $J(\hat{\mathbf{z}})$ are real) in the sense that it is a block diagonal matrix with block size at most $2\times 2$ and $Q$ is a unitary matrix. Then, we have $J(\hat{\mathbf{z}})+\lambda I=Q(U+\lambda I)Q^{-1}$. In the convex-concave setting, $J(\hat{\mathbf{z}})$ is positive semidefinite and thus $J(\hat{\mathbf{z}})+\lambda I$ is positive definite.
@@
$\Delta\mathbf{z}(\lambda)\triangleq-(J(\hat{\mathbf{z}})+\lambda I)^{-1}F(\hat{\mathbf{z}})=-Q(U+\lambda I)^{-1}Q^{-1}F(\hat{\mathbf{z}}),$ (4.10)
@@
By differentiating the equation $(J(\hat{\mathbf{z}})+\lambda I)\Delta\mathbf{z}(\lambda)=-F(\hat{\mathbf{z}})$, we have
@@
Putting these pieces together yields the desired expressions of $\psi^{\prime}(\lambda)$ and $\psi^{\prime\prime}(\lambda)$. Since $F(\hat{\mathbf{z}})\neq\textbf{0}$ and $J(\hat{\mathbf{z}})$ is positive semidefinite, we have $\psi^{\prime}(\lambda)<0$ and $\psi^{\prime\prime}(\lambda)\geq 0$ when $\lambda>0$. This completes the proof. $\Box$
@@
$\phi^{\prime\prime}(\lambda)$
$=$ $\tfrac{\|\Delta\mathbf{z}(\lambda)\|^{2}(\|(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda)\|^{2}+2\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-2}\Delta\mathbf{z}(\lambda))-(\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda))^{2}}{\|\Delta\mathbf{z}(\lambda)\|^{3}}$
@@
$\geq$ 0,when $0,\quad\textnormal{when }\lambda>0,$
where the first inequality holds true since $J(\hat{\mathbf{z}})$ is positive semidefinite, $F(\hat{\mathbf{z}})\neq\textbf{0}$ and $\lambda>0$, and the second inequality holds true because of the the Cauchy-Schwartz inequality. This completes the proof. $\Box$
The required solution is the unique positive root to Eq. (4.11) since $\phi(\lambda)$ is decreasing. Let $\lambda^{0}>0$ be given with $\phi(\lambda^{0})>0$, we first perform one Schur decomposition of $J(\hat{\mathbf{z}})$:
@@
A typical iteration of the unit-stepsize Newton method for finding such root replaces the current iterate $\lambda^{j}>0$ with the improved estimate $\lambda^{j+1}$ for which
@@
The value of $\phi(\lambda^{j})$ can be obtained by solving $\Delta\mathbf{z}(\lambda^{j})=-Q(U+\lambda^{j}I)^{-1}Q^{-1}F(\hat{\mathbf{z}})$, and that of $\phi^{\prime}(\lambda^{j})$ can be obtained using Proposition 4.5 once $\Delta\mathbf{z}(\lambda^{j})$ is available (solving a linear system is required). This implies that $\phi(\lambda^{j})$ and $\phi^{\prime}(\lambda^{j})$ can be computed by two calls to a linear system solver in a quasi-upper-triangle system.
In terms of complexity, the cost of a single Schur decomposition is $O((m+n)^{\omega})$ where $\omega\approx 2.3728$ is the matrix multiplication constant and the cost of one Newton iteration is $O((m+n)^{2})$. In terms of convergence, the unit-stepsize Newton method by itself is not a reliable method and the iterates do not converge in general. However, the convexity of $\phi$ established in Proposition 4.5 has useful properties.
@@
Suppose that the iterates $\{\lambda^{j}\}_{j\geq 0}$ are generated by a unit-stepsize Newton method with $\lambda^{0}>0$ with $\phi(\lambda^{0})\geq 0$. Then, we have $\lambda^{j}>0$ and $\phi(\lambda^{j})\geq 0$. As a consequence, the iterates converge monotonically towards the unique solution $\lambda^{\star}$. The convergence rate is globally $Q$-linear with a factor at least $1-\phi^{\prime}(\lambda^{\star})/\phi^{\prime}(\lambda^{0})$ and is ultimately $Q$-quadratic.
@@
$10\leq 1-\tfrac{\phi^{\prime}(\tilde{\lambda})}{\phi^{\prime}(\lambda^{j})}\leq 1-\tfrac{\phi^{\prime}(\lambda^{\star})}{\phi^{\prime}(\lambda^{0})}<$ 1 .
Putting everything together implies that the convergence rate is globally $Q$-linear with a factor at least $1-\phi^{\prime}(\lambda^{\star})/\phi^{\prime}(\lambda^{0})$. The asymptotic $Q$-quadratic convergence of the Newton iteration follows because the Jacobian of $\phi$ is nonsingular at the unique solution $\lambda^{\star}$ (i.e., $\phi^{\prime}(\lambda^{\star})\neq 0$). This completes the proof. $\Box$
@@
Our approach is inspired by Adil et al. (2022), who reformulated the subproblem in their algorithm as the nonlinear equality $\lambda=6\rho\|\Delta\mathbf{z}(\lambda)\|$ and proposed a bisection method as a subroutine. In total, this requires a single Schur decomposition and $O(\log(1/\epsilon))$ calls to a linear system solver for a quasi-upper-triangular system. Our key finding is that the Newton method is better than the bisection method for solving the nonlinear equation $\lambda=6\rho\|\Delta\mathbf{z}(\lambda)\|$. Indeed, the simple reformulation in Eq. (4.11) and the unit-stepsize implementation achieve global linear and local quadratic convergence guarantees. Thus, by comparison, the proposed subroutine only requires a single Schur decomposition and $O(\log\log(1/\epsilon))$ calls to a linear system solver for a quasi-upper-triangular system.
@@
We note also that Huang et al. (2022) reformulated the subproblem in their algorithm as the two-dimensional nonlinear equation problem (see Huang et al. (2022, Eq.(16))) and proposed the Newton method as a subroutine. However, no theoretical guarantee was provided, even in a local sense; indeed, the proposed nonlinear equation is complex and it is not clear if it satisfies essential properties (e.g., the nonsingularity of Jacobian). Our approach is similar in spirit to Conn et al. (2000); they consider the choice of $\phi\triangleq 1/\sqrt{\psi(\lambda)}-6\rho/\lambda$ and prove that it is concave in optimization setting where the second-order information is a symmetric matrix. Their analysis is unfortunately not valid here since $J(\hat{\mathbf{z}})$ is, in general, not symmetric.
@@ 4.4 Finite-sum min-max optimization
We give concrete examples to clarify the ways to construct the inexact Jacobian such that Condition 4.1 holds true. The key ingredient is random sampling which can significantly reduce the computational cost in an optimization setting (Xu et al., 2020) and we show that such technique can be employed for solving finite-sum min-max optimization problems in Eq. (4.2) and (4.3).
@@
$J(\mathbf{z})=\tfrac{1}{N|\mathcal{S}|}\sum_{i\in\mathcal{S}}\tfrac{1}{p_{i}}DF_{i}(\mathbf{z}),\quad\textnormal{for }DF_{i}(\mathbf{z})=\begin{bmatrix}\nabla^{2}_{\textbf{xx}}f_{i}(\mathbf{x},\mathbf{y})&\nabla^{2}_{\textbf{xy}}f_{i}(\mathbf{x},\mathbf{y})\\ -\nabla^{2}_{\textbf{xy}}f_{i}(\mathbf{x},\mathbf{y})&-\nabla^{2}_{\textbf{yy}}f_{i}(\mathbf{x},\mathbf{y})\end{bmatrix}.$ (4.13)
@@
$\sup_{\mathbf{z}}\|DF_{i}(\mathbf{z})\|\leq B_{i},\quad\textnormal{for all }i\in\{1,2,\ldots,N\},$ (4.14)
and let $B_{\max}=\max_{1\leq i\leq N}B_{i}$. For the uniform sampling (i.e., $p_{i}=\frac{1}{N}$), we summarize the sample complexity results in the following lemma. The proof is omitted for brevity.
@@
Suppose that Eq. (4.14) holds true and let $B_{\max}$ and $10<\tau,\delta<$ 1 be defined properly. A uniform sampling with or without replacement is performed to form the subsampled Jacobian; indeed, $J(\mathbf{z})$ is constructed using Eq. (4.13) with $p_{i}=\frac{1}{n}$ and the sample size satisfies
@@
Lemma 4.9 shows that the inexact Jacobian satisfies Condition 4.1 with probability $1-\delta$ under certain $\tau$ and $\kappa_{J}=B_{\max}$ if it is constructed using the uniform sampling and the size $|\mathcal{S}|=\Omega(\frac{B_{\max}^{2}}{\tau^{2}}\log(\tfrac{m+n}{\delta}))$. Indeed, the first inequality holds true with probability $1-\delta$ since $\textnormal{Prob}(\|J(\mathbf{z})-DF(\mathbf{z})\|\leq\tau)\geq 1-\delta$, and the second inequality holds true since $\kappa_{H}=B_{\max}$ (this is a deterministic statement).
In the special finite-sum setting with Eq. (4.3), we can construct a more “informative” distribution of sampling $\xi\in\{1,2,\ldots,N\}$ as opposed to simplest uniform sampling. It is advantageous to bias the probability distribution towards carefully picking indices corresponding to those relevant $f_{i}$’s in forming the Jacobian. However, constructing inexact Hessian and corresponding sample complexity guarantee from Xu et al. (2020, Section 3.1) requires $\nabla^{2}f_{i}$ to be rank-one, which is not valid here. To address this issue, we avail ourselves of the operator-Bernstein inequality (Gross and Nesme, 2010).
The Jacobian of $F$ can be rewritten as $DF(\mathbf{z})=\tfrac{1}{N}\sum_{i=1}^{N}\Lambda_{i}DF_{i}(\mathbf{a}_{i}^{\top}\mathbf{x},\mathbf{b}_{i}^{\top}\mathbf{y})\Lambda_{i}^{\top}$ where $\Lambda_{i}=\begin{bmatrix}\mathbf{a}_{i}&\\ &\mathbf{b}_{i}\end{bmatrix}\in\mathbb{R}^{(m+n)\times 2}$ and $DF_{i}(x,y)\in\mathbb{R}^{2\times 2}$. Then, the resulting compact form is $DF(\mathbf{z})=\Lambda^{\top}\Sigma\Lambda$ where
@@
We suppose
$\sup_{(\mathbf{x},\mathbf{y})}\|DF_{i}(\mathbf{a}_{i}^{\top}\mathbf{x},\mathbf{b}_{i}^{\top}\mathbf{y})\|(\|\mathbf{a}_{i}\|^{2}+\|\mathbf{b}_{i}\|^{2})\leq B_{i},\quad\textnormal{for all }i\in\{1,2,\ldots,N\},$ (4.16)
and let $B_{\textnormal{avg}}=\frac{1}{N}\sum_{i=1}^{N}B_{i}$. For the uniform sampling with the particular nonuniform distribution given by
$p_{i}=\tfrac{\|DF_{i}(\mathbf{a}_{i}^{\top}\mathbf{x},\mathbf{b}_{i}^{\top}\mathbf{y})\|(\|\mathbf{a}_{i}\|^{2}+\|\mathbf{b}_{i}\|^{2})}{\sum_{i=1}^{N}\|DF_{i}(\mathbf{a}_{i}^{\top}\mathbf{x},\mathbf{b}_{i}^{\top}\mathbf{y})\|(\|\mathbf{a}_{i}\|^{2}+\|\mathbf{b}_{i}\|^{2})}.$ (4.17)
@@
Suppose that Eq. (4.16) holds true and let $B_{\textnormal{avg}}$ and $10<\tau,\delta<$ 1 be defined properly. A nonuniform sampling is performed to form the subsampled Jacobian; indeed, $J(\mathbf{z})$ is constructed using Eq. (4.13) with $p_{i}>0$ in Eq. (4.17) and the sample size satisfies
@@
Compared to Lemma 4.9, the computation of sampling probability in Lemma 4.11 requires going through the whole dataset and the cost is $O((m+n)N)$. Nonetheless, the computational savings with smaller sample size dominates such extra cost of computing the sampling probability in optimization setting (Xu et al., 2016). In particular, the sample size from Lemma 4.11 is smaller as $B_{\textnormal{avg}}\ll B_{\max}$ which occurs if one $B_{i}$ is much larger than the others. In addition, the sample size is proportional to the log of the failure probability in Lemma 4.9 and 4.11, allowing the use of a very small failure probability without increasing the sample size significantly.
@@
     STEP 3: Compute an inexact solution $\Delta\mathbf{z}_{k}$ of the following subproblem
@@
     STEP 4: Compute $\lambda_{k+1}>0$ such that $\frac{1}{30}\leq\lambda_{k+1}\rho\|\Delta\mathbf{z}_{k}\|\leq\frac{1}{14}$.
     STEP 5: Compute $\mathbf{z}_{k+1}=\hat{\mathbf{z}}_{k}+\Delta\mathbf{z}_{k}$.
     STEP 6: Compute $\hat{\mathbf{z}}_{k+1}=\hat{\mathbf{z}}_{k}-\lambda_{k+1}F(\mathbf{z}_{k+1})$.
@@
Combining Algorithm 2 and these random sampling strategies gives the first class of subsampled Newton methods for solving finite-sum min-max optimization problems. We provide the scheme in Algorithm 3 and present the iteration complexity and total complexity guarantees in the high-probability sense.
@@
Suppose that Assumption 2.4 and 2.5 hold. Then, the iterates generated by Algorithm 3 are bounded and Algorithm 3 achieves an $\epsilon$-global saddle point solution within $O(\epsilon^{-2/3})$ iterations with the probability at least $1-\delta$. The total complexity bound of inexact methods is $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the matrix multiplication constant.

### v5

@@ 4.3 Inexact subproblem solving and complexity analysis
We now clarify how to obtain an inexact solution of the nonlinear equation problem in Eq. (4.1) such that Condition 4.2 holds. In particular, we present a new subroutine that solves each subproblem using a single Schur decomposition and only $O(\log\log(1/\epsilon))$ calls to a linear system solver in a quasi-upper-triangular system. This gives a total complexity of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ which outperforms that of $O((m+n)^{\omega}\epsilon^{-2/3}\log\log(1/\epsilon))$ achieved by the best known line-search-based method.
For simplicity, we omit the subscript and rewrite the nonlinear equation problem in Eq. (4.1) as
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\phi^{\prime\prime}(\lambda)$ $=$ $\tfrac{\|\Delta\mathbf{z}(\lambda)\|^{2}(\|(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda)\|^{2}+2\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-2}\Delta\mathbf{z}(\lambda))-(\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda))^{2}}{\|\Delta\mathbf{z}(\lambda)\|^{3}}$
@@
$\geq$ $0,\quad\textnormal{when }\lambda>0,$
[…]
@@
A typical iteration of the Newton method for finding such root replaces the current iterate $\lambda^{j}>0$ with the improved estimate $\lambda^{j+1}$ for which
@@
[…]
@@
Suppose that the iterates $\{\lambda^{j}\}_{j\geq 0}$ are generated by the Newton method with $\lambda^{0}>0$ with $\phi(\lambda^{0})\geq 0$. Then, we have $\lambda^{j}>0$ and $\phi(\lambda^{j})\geq 0$. As a consequence, the iterates converge monotonically towards the unique solution $\lambda^{\star}$. The convergence rate is globally $Q$-linear with a factor at least $1-\phi^{\prime}(\lambda^{\star})/\phi^{\prime}(\lambda^{0})$ and is ultimately $Q$-quadratic.
@@
$0\leq 1-\tfrac{\phi^{\prime}(\tilde{\lambda})}{\phi^{\prime}(\lambda^{j})}\leq 1-\tfrac{\phi^{\prime}(\lambda^{\star})}{\phi^{\prime}(\lambda^{0})}<1.$
Putting everything together then implies that the convergence rate is globally $Q$-linear with a factor at least $1-\phi^{\prime}(\lambda^{\star})/\phi^{\prime}(\lambda^{0})$. The asymptotic $Q$-quadratic convergence of the Newton iteration follows because the Jacobian of $\phi$ is nonsingular at the unique solution $\lambda^{\star}$ (i.e., $\phi^{\prime}(\lambda^{\star})\neq 0$). $\Box$
@@
Our reformulation closely follows Adil et al. (2022) who proposed a bisection method as a solver and proved a complexity bound of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the matrix multiplication constant (Demmel et al., 2007). We also note that Huang et al. (2022) reformulated the same subproblem as a two-dimensional nonlinear equation (see Huang et al. (2022, Eq.(16))) and proposed a damped Newton method as a solver but without any complexity bound guarantee. Our finding is that the Newton method outperforms bisection and damped Newton in that it leverages the specific structure of Eq. (4.9) to yield not only an improved complexity bound but a practical implementation.
@@
Our analysis in Theorem 4.6 is related to Cartis et al. (2011, Theorem 6.3). The key difference lies in how we adapt their techniques to settings involving asymmetric Jacobian matrices. Specifically,  Cartis et al. (2011) analyze the function $\phi(\lambda)=1/\sqrt{\psi(\lambda)}-6\rho/\lambda$ and show that it is concave when the second-order information is symmetric. However, this function will not be concave in the presence of asymmetry. We address this issue by proposing a different function, $\phi(\lambda)=\sqrt{\psi(\lambda)}-\lambda/(6\rho)$, which we can prove remains concave even when the Jacobian matrix is asymmetric. This change is crucial because it allows us to retain the theoretical convergence guarantees of the Newton method without damping, even in the non-symmetric case.
@@ 4.4 Finite-sum min-max optimization
We give concrete examples to clarify the ways to construct the inexact Jacobian such that Condition 4.1 holds true. The key ingredient is random sampling which can significantly reduce the computational cost in a minimization setting Xu et al. (2020) and we show that such technique can be employed for solving finite-sum min-max optimization problems in Eq. (4.2) and (4.3).
@@
[…]
@@
$\sup_{\mathbf{z}\in\mathbb{R}^{m+n}}\|DF_{i}(\mathbf{z})\|\leq B_{i},\quad\textnormal{for all }i\in\{1,2,\ldots,N\},$ (4.14)
and let $B_{\max}=\max_{1\leq i\leq N}B_{i}$. The sample complexity results for the uniform sampling (i.e., $p_{i}=\frac{1}{N}$) are a consequence of Xu et al. (2020, Lemma 16) and we omit the proof for the brevity.
@@
Suppose that Eq. (4.14) holds true and let $B_{\max}$ and $0<\tau,\delta<1$ be defined properly. A uniform sampling with or without replacement is performed to form the subsampled Jacobian; indeed, $J(\mathbf{z})$ is constructed using Eq. (4.13) with $p_{i}=\frac{1}{n}$ and the sample size satisfies
@@
Lemma 4.9 shows that the inexact Jacobian satisfies Condition 4.1 with probability $1-\delta$ under certain choice of $\tau$ and $\kappa_{J}=B_{\max}$ if it is constructed using the uniform sampling and the size $|\mathcal{S}|=\Omega(\frac{B_{\max}^{2}}{\tau^{2}}\log(\tfrac{m+n}{\delta}))$. Indeed, the first inequality holds true with probability $1-\delta$ since $\textnormal{Prob}(\|J(\mathbf{z})-DF(\mathbf{z})\|\leq\tau)\geq 1-\delta$, and the second inequality holds true since $\kappa_{H}=B_{\max}$ (this is deterministic).
In the finite-sum setting with Eq. (4.3), we construct a more “informative” distribution of sampling $\xi\in\{1,2,\ldots,N\}$ as opposed to simplest uniform sampling. It is advantageous to bias the probability distribution towards carefully picking indices corresponding to those relevant $f_{i}$’s in forming the Jacobian. However, constructing inexact Hessian and corresponding sample complexity guarantee from Xu et al. (2020, Section 3.1) in a minimization setting requires $\nabla^{2}f_{i}$ to be rank-one, which is not valid here. To address this, we avail ourselves of the operator-Bernstein inequality (Gross and Nesme, 2010).
We write $DF(\mathbf{z})=\tfrac{1}{N}\sum_{i=1}^{N}\Lambda_{i}DF_{i}(\mathbf{a}_{i}^{\top}\mathbf{x},\mathbf{b}_{i}^{\top}\mathbf{y})\Lambda_{i}^{\top}$ where $\Lambda_{i}=\begin{bmatrix}\mathbf{a}_{i}&\\ &\mathbf{b}_{i}\end{bmatrix}\in\mathbb{R}^{(m+n)\times 2}$ and $DF_{i}(x,y)\in\mathbb{R}^{2\times 2}$. Then, the resulting compact form is $DF(\mathbf{z})=\Lambda^{\top}\Sigma\Lambda$ where
@@
We assume that $B_{\textnormal{avg}}=\frac{1}{N}\sum_{i=1}^{N}B_{i}$ and
$\sup_{(\mathbf{x},\mathbf{y})}\|DF_{i}(\mathbf{a}_{i}^{\top}\mathbf{x},\mathbf{b}_{i}^{\top}\mathbf{y})\|(\|\mathbf{a}_{i}\|^{2}+\|\mathbf{b}_{i}\|^{2})\leq B_{i},\quad\textnormal{for all }i\in\{1,2,\ldots,N\}.$ (4.16)
For the uniform sampling with the particular nonuniform distribution given by
$p_{i}=\tfrac{\|DF_{i}(\mathbf{a}_{i}^{\top}\mathbf{x},\mathbf{b}_{i}^{\top}\mathbf{y})\|(\|\mathbf{a}_{i}\|^{2}+\|\mathbf{b}_{i}\|^{2})}{\sum_{i=1}^{N}\|DF_{i}(\mathbf{a}_{i}^{\top}\mathbf{x},\mathbf{b}_{i}^{\top}\mathbf{y})\|(\|\mathbf{a}_{i}\|^{2}+\|\mathbf{b}_{i}\|^{2})},$ (4.17)
@@
Suppose that Eq. (4.16) holds and let $B_{\textnormal{avg}}$ and $0<\tau,\delta<1$ be defined properly. Then, the nonuniform sampling is performed to form the subsampled Jacobian; indeed, $J(\mathbf{z})$ is constructed using Eq. (4.13) with $p_{i}>0$ in Eq. (4.17) and the sample size satisfies
@@
Compared to Lemma 4.9, the computation of sampling probability in Lemma 4.11 requires going through the whole dataset and the cost is $O((m+n)N)$. Nonetheless, the computational savings with smaller sample size dominates such extra cost of computing the sampling probability in a minimization setting (Xu et al., 2016). In particular, the sample size from Lemma 4.11 is smaller as $B_{\textnormal{avg}}\ll B_{\max}$ which occurs if one $B_{i}$ is much larger than the others. In addition, the sample size is proportional to the log of the failure probability in Lemma 4.9 and 4.11, allowing the use of a very small failure probability without increasing the sample size significantly.
@@
  STEP 2: Compute an inexact solution $\Delta\mathbf{z}_{k}$ of the following subproblem
@@
  STEP 4: Compute $\mathbf{z}_{k+1}=\hat{\mathbf{z}}_{k}+\Delta\mathbf{z}_{k}$.
@@
Combining Algorithm 2 and these random sampling strategies gives the first class of subsampled Newton methods for solving finite-sum min-max optimization problems. We summarize the scheme in Algorithm 3 and present the iteration complexity and total complexity guarantees with high-probability in the following theorem.
@@
Suppose that Assumptions 2.4 and 2.5 hold. Then, the iterates generated by Algorithm 3 are bounded and Algorithm 3 achieves an $\epsilon$-global saddle point solution within $O(\epsilon^{-2/3})$ iterations with the probability at least $1-\delta$. As a consequence, the high-probability complexity bound is $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the multiplication constant.
Remark 4.14
In the context of finite-sum optimization, Arjevani and Shamir (2017) investigated lower bounds for second-order methods, where the complexity depends on the number of samples, the condition number, and the desired tolerance. They identified a gap between upper and lower bounds for subsampled Newton methods, indicating that simple sampling strategies are insufficient and that variance reduction or averaging could be essential for further improvement. Recent advances in stochastic second-order optimization methods with variance reduction further indicate the importance of combining structure and noise control (Zhou et al., 2019; Na et al., 2023). However, the lower bound on the oracle complexity of second-order methods for finite-sum minimax optimization remains an open question. This is beyond the scope of our current work but represents an important direction for future research.